OpenStudy (anonymous):

The molar mass (molecular weight) of NaOH is 40.0. How many grams of NaOH will be required to prepare 0.01 NaOH solution A) 40.0 g B) 4.0 g C) 0.40 g D) 0.04 g

OpenStudy (anonymous):

Your making a 0.01 Molar NaOH solution?

OpenStudy (anonymous):

If so it would be C, as Mass = Molar Mass * Moles So 40 * 0.01 = 0.4 :)

OpenStudy (anonymous):

no but the answer is .04!

OpenStudy (anonymous):

that is D

thomaster (thomaster):

@C9H8O4 is right. C is the correct answer. 1 mol = 40 gram 0.01 mol = 0.01*40 = 0.4 gram.

OpenStudy (anonymous):

oh ok.